# inkscapelatex=false

### v1

@@ A Riemannian metric $g$ on $(\mathbb{R}_{0}^{n+1})^{\ell}$ may be defined as follows:
for each $X_{\mathbf{x}}=(X_{{\rm x}_{1}},\ldots,X_{{\rm x}_{\ell}}),Y_{\mathbf{x}}=(Y_{{\rm x}_{1}},\ldots,Y_{{\rm x}_{\ell}})\in T_{\mathbf{x}}(\mathbb{R}_{0}^{n+1})^{\ell}$ and $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in(\mathbb{R}_{0}^{n+1})^{\ell}$. The orthogonal decomposition induced by $g$ is denoted by:
$T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}=T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}\oplus T_{\mathbf{y}}^{\perp}(\mathcal{E}_{P}^{n})^{\ell},\quad X_{\mathbf{y}}=X_{\mathbf{y}}^{\parallel}+X_{\mathbf{y}}^{\perp},\qquad\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell},$
@@
$T_{\mathbf{y}}^{\perp}(\mathcal{E}_{P}^{n})^{\ell}=\{X_{\mathbf{y}}\in T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}\mid\langle X_{\mathbf{y}},Y_{\mathbf{y}}\rangle_{g(\mathbf{y})}=0,~{}\forall\,Y_{\mathbf{y}}\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}\},\quad\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}.$
@@
The orthogonal projection is given by $\boldsymbol{\pi}^{\parallel}=T\boldsymbol{\pi}_{P}|_{(\mathcal{E}_{P}^{n})^{\ell}}$.
@@
It is enough to prove that, for each $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{P}^{n})^{\ell}$ and $X_{\mathbf{y}}=(X_{{\rm y}_{1}},\ldots,X_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}$, we have that $X_{\mathbf{y}}-T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{\mathbf{y}}\in T_{\mathbf{y}}^{\top}(\mathcal{E}_{P}^{n})^{\ell}$, i.e., that $\langle X_{\mathbf{y}}-T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{\mathbf{y}},Y_{\mathbf{y}}\rangle_{g(\mathbf{y})}=0$ for each $Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. By using (3) and (9), this latter condition is clearly satisfied:
@@
where we have used Remark 2.1. ∎
@@ 3.2. Gradient vector field
Let us show that the transformer dynamics is a gradient vector field on the manifold $(\mathcal{E}_{P}^{n})^{\ell}$ equipped with the Riemannian metric $g_{P}=\imath_{P}^{*}g$. For simplicity, we introduce the following vector fields corresponding to (8) before and after projecting to the ellipsoid, respectively:
@@
$Y_{P}:\mathcal{E}^{n}(P)^{\ell}\to T\mathcal{E}^{n}(P)^{\ell},\quad\mathbf{y}\mapsto Y_{P}(\mathbf{y})=T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{P}(\mathbf{y})=\begin{pmatrix}T_{{\rm y}_{1}}\pi_{P}\cdot\left(\sum_{j=1}^{\ell}\alpha_{1j}(\mathbf{y})\,{\rm y}_{j}\right)\\ \vdots\\ T_{{\rm y}_{\ell}}\pi_{P}\cdot\left(\sum_{j=1}^{\ell}\alpha_{\ell j}(\mathbf{y})\,{\rm y}_{j}\right)\end{pmatrix}.$
@@
From this, (1) and (9), we conclude:
${\rm grad}_{g}\,V(\mathbf{x})$ $=-\begin{pmatrix}Z_{1}^{-1}(\mathbf{x})\,P^{-1}\,&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&Z_{\ell}^{-1}(\mathbf{x})\,P^{-1}\,\end{pmatrix}\begin{pmatrix}\sum_{j=1}^{\ell}\exp({\rm x}_{1}^{\top}\,P\,{\rm x}_{j})\,P\,{\rm x}_{j}\\ \vdots\\ \sum_{i=j}^{\ell}\exp({\rm x}_{\ell}^{\top}\,P\,{\rm x}_{j})\,P\,{\rm x}_{j}\end{pmatrix}=-X_{P}(\mathbf{x}).$
@@
where we used Lemma 3.1 and the equality $T_{\mathbf{y}}\imath_{P}\cdot Y_{P}(\mathbf{y})=Y_{P}(\mathbf{y})$, which follows from regarding $Y_{P}(\mathbf{y})$ both as an element of $T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}$ and $T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. ∎
@@ 3.3. Stability analysis
Having established that (8) is a gradient vector field, it is natural to use the potential $V_{P}:(\mathcal{E}_{P}^{n})^{\ell}\to\mathbb{R}$ as a Lyapunov function to study the asymptotic behavior of the tokens.
Lemma 3.3.
The trajectories of the system (8) converge to the set:
@@
Let $\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}$ and $Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. Recall that the formal time derivative (at $t=0$) of the potential $V_{P}$ is the map $\dot{V}_{P}=\mathbf{d}V_{P}(Y_{P}):(\mathcal{E}_{P}^{n})^{\ell}\to\mathbb{R}$. From Theorem 3.1, we obtain:
@@
Every trajectory of the system (8) converges to an equilibrium.
@@
Recall that the potential $V_{P}$ satisfies the Łojasiewicz inequality if $|V_{P}|\leq\lambda\,|{\rm grad}_{g_{P}}\,V_{P}|_{h}$ for some $\lambda>0$. A sufficient condition for the Łojasiewicz inequality to hold is that $((\mathcal{E}_{P}^{n})^{\ell},g_{P}=\imath^{*}g)$ is a real analytic, Riemannian manifold and the potential is real analytic, i.e., $V_{P}\in C^{\omega}((\mathcal{E}_{P}^{n})^{\ell})$. It is clear that these two conditions are satisfied since $(\mathcal{E}_{P}^{n})^{\ell}$ is a real analytic submanifold of $\mathbb{R}^{n+1}$ and $Z_{i},\alpha_{ij}\in C^{\omega}((\mathbb{R}_{0}^{n+1})^{\ell}),\mathbb{R}^{+})$ for each $1\leq i,j\leq\ell$, which ensures that both the Riemannian metric $g_{P}$ and the potential $V_{P}$ are real analytic.
On the other hand, $(\mathcal{E}_{P}^{n})^{\ell}\subset(\mathbb{R}_{0}^{n+1})^{\ell}$ is compact, whence the set of $\omega$-limit points of (8) is non-empty. The Łojasiewicz inequality thus ensures that every trajectory converges to a point $\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}$. From Lemma 3.3, we know that $\mathbf{y}$ is an equilibrium of (8) since $\dot{\mathbf{y}}=-{\rm grad}_{g_{P}}\,V_{P}(\mathbf{y})=0$. ∎
@@
4. Tokens evolving on an hemisphere
The conditions assumed in the previous section are quite restrictive and hardly encountered in real transformers: only one head and $P$ time invariant, symmetric, and positive definite. For that reason, in this section we consider the particular case of (7) with identity value matrices, $U_{\eta}(t)=\mathbb{I}_{n+1}$ and $P_{\eta}(t)$ uniformly continuous and bounded, but not necessarily symmetric nor positive definite, for each head $1\leq\eta\leq h$. The matrix $W$ is assumed to be symmetric and positive-definite. Under these assumptions, (7) becomes:
(11) $\dot{\rm y}_{i}=T_{{\rm y}_{i}}\pi_{W}\cdot\left(\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y})\,{\rm y}_{j}\right)=\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y})\left({\rm y}_{j}-{\rm y}_{i}^{\top}\,W\,{\rm y}_{j}\,{\rm y}_{i}\right),$
for each $1\leq i\leq\ell$, $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{W}^{n})^{\ell}$, where:
$\alpha_{ij}^{\eta}(t,\mathbf{y})=\frac{1}{Z_{i}^{\eta}(t,\mathbf{y})}\,\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}),\qquad Z_{i}^{\eta}(t,\mathbf{y})=\sqrt{n+1}\,\sum_{j=1}^{\ell}\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}).$
In the proof of Theorem 4.1 below, a nonsmooth candidate for Lyapunov function will be introduced. In order to handle this situation, we briefly recall how to compute the Dini derivative of a function defined through a maximum (cf. [30, §2.3]).
@@ 4. Tokens evolving on an hemisphere
As a particular case, let $\{f_{i}:[0,\infty[\to\mathbb{R}\mid i\in I\}$ be a family of continuously differentiable functions, and consider its maximum:
$f:[0,\infty[\to\mathbb{R},\quad t\mapsto f(t)=\max_{i\in I}f_{i}(t).$
For each $t\geq 0$, Danskin’s theorem yields the upper Dini derivative of $f$(cf. [30, Lemma 2.2]):
@@
where $\mathcal{I}(t)=\{i\in I\mid f(t)=f_{i}(t)\}$.
In addition, let us introduce the following two technical lemmas.
@@
If there exists $b>0$ such that $\max_{1\leq\eta\leq h}\sup_{t\geq 0}\|P_{\eta}(t)\|\leq b$, then there exist $c_{1},c_{2}>0$ such that $c_{1}\leq\alpha_{ij}^{\eta}(t,\mathbf{y})\leq c_{2}$ for each $t\geq 0$, $\mathbf{y}\in(\mathcal{E}_{W}^{n})^{\ell}$, $1\leq\eta\leq h$ and $1\leq i,j\leq\ell$.
@@
By compactness of $\mathcal{E}_{W}^{n}$, there exists $K>0$ such that $|{\rm y}|\leq K$ for each ${\rm y}\in\mathcal{E}_{W}^{n}$. Therefore, we have $|{\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j}|\leq K^{2}\,\|P_{\eta}(t)\|\leq K^{2}\,b$ and, thus, $\exp(-K^{2}\,b)\leq\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{j})\leq\exp(K^{2}\,b)$. Moreover, we also have $\exp(-K^{2}\,b)\leq\sum_{k=1}^{\ell}\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{k})\leq\ell\,\exp(K^{2}\,b)$, which leads to:
$\alpha_{ij}^{\eta}(t,\mathbf{y})=\frac{\exp({\rm y}_{i}^{\top}\,P_{\eta}(t){\rm y}_{j})}{\sqrt{n+1}\,\sum_{k=1}^{i}\exp({\rm y}_{i}^{\top}\,P_{\eta}(t)\,{\rm y}_{k})}\geq\frac{\exp(-K^{2}\,b)}{\sqrt{n+1}\,\ell\,\exp(K^{2}\,b)}\geq\frac{1}{\sqrt{n+1}\,\ell\,\exp(2\,K^{2}\,b)}.$
@@
Let $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell}):[0,\infty[\to\mathcal{E}_{W}^{n}$ be a solution of (11) and $v\in\mathcal{E}_{W}^{n}$. If there exists $b>0$ such that $\sup_{t\geq 0}\|P_{\eta}(t)\|\leq b$ and $P_{\eta}$ is uniformly continuous on [0,[[0$\infty$[[ 0 , ∞ [ for each $1\leq\eta\leq h$, then the following functions:
$f_{i}:[0,\infty[\to\mathbb{R},\quad t\mapsto f_{i}(t)=\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y}(t))\left(v^{\top}\,{\rm y}_{j}(t)-{\rm y}_{i}(t)^{\top}\,W\,{\rm y}_{j}(t)\,v^{\top}\,{\rm y}_{i}(t)\right),$
@@
$g_{j}:[0,\infty[\to\mathbb{R},\quad$ $t\mapsto g_{j}(t)=v^{\top}\,{\rm y}_{j}(t)-{\rm y}_{i}(t)^{\top}\,W\,{\rm y}_{j}(t)\,v^{\top}\,{\rm y}_{i}(t),$
$g_{j}^{\eta}:[0,\infty[\to\mathbb{R},\quad$ $t\mapsto g_{j}^{\eta}(t)=\alpha_{ij}^{\eta}(t,\mathbf{y}(t)),$
for each $1\leq\eta\leq h$ and $1\leq j\leq\ell$, are bounded: $g_{j}$ due to tokens evolving on the ellipsoid, and $g_{j}^{\eta}$ due to Lemma 4.1. Recall that the addition and multiplication of bounded and uniformly continuous functions results in uniformly continuous functions. Thus, it is enough to prove that $g_{j}$ and $g_{j}^{\eta}$ are uniformly continuous to conclude that $f_{i}$ is uniformly continuous:
@@
    The derivative of $g_{j}$ is bounded on [0,[[0$\infty$[[ 0 , ∞ [ since the tokens evolve on the ellipsoid and their dynamics is given by (11). Note that $\sup_{t\geq 0}|\dot{\rm y}_{j}(t)|<\infty$ for each $1\leq j\leq\ell$ thanks to Lemma 4.1. Hence, $g_{j}$ is uniformly continuous on [0,[[0$\infty$[[ 0 , ∞ [.
@@
    Given that the tokens evolve on the ellipsoid and every $P_{\eta}$ is bounded on [0,[[0$\infty$[[ 0 , ∞ [, we can ensure the existence of $K>0$ such that $\max_{1\leq\eta\leq h}\sup_{t\geq 0}|{\rm y}_{i}(t)^{\top}\,P_{\eta}(t)\,{\rm y}_{j}(t)|\leq K$. Moreover, $\exp|_{[-K,K]}:[-K,K]\to[\exp(-K),\exp(K)]$ is uniformly continuous, as it is defined on a compact. Hence, $\exp({\rm y}_{i}(\cdot)\,P_{\eta}(\cdot)\,{\rm y}_{j}(\cdot)):[0,\infty[\to[\exp(-K),\exp(K)]$ is uniformly continuous, as the composition of uniformly continuous functions is uniformly continuous. In particular, $Z_{i}^{\eta}(\cdot,\mathbf{y}(\cdot)):[0,\infty[\to[\ell\,\exp(-K),\ell\,\exp(K)]$ is uniformly continuous. By gathering all, we conclude that $g_{j}^{\eta}$ is uniformly continuous on [0,[[0$\infty$[[ 0 , ∞ [.
We now use the previous results to prove attractivity of the consensus set, introduced in the next result, provided that the initial position of the tokens is some open hemisphere of the ellipsoid.
@@
If there exists $b>0$ such that $\sup_{t\geq 0}\|P_{\eta}(t)\|\leq b$ and $P_{\eta}$ is uniformly continuous on [0,[[0$\infty$[[ 0 , ∞ [ for each $1\leq\eta\leq h$, then the consensus set on the hemisphere:
@@
is attractive for (11) with domain of attraction $\mathcal{H}^{+}(v)^{\ell}$.
@@
For each $t\geq 0$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in\mathcal{H}^{+}(v)$, note that:

### v2

@@ A Riemannian metric $g$ on $(\mathbb{R}_{0}^{n+1})^{\ell}$ may be defined as follows:
for each $X_{\mathbf{x}}=(X_{{\rm x}_{1}},\ldots,X_{{\rm x}_{\ell}}),Y_{\mathbf{x}}=(Y_{{\rm x}_{1}},\ldots,Y_{{\rm x}_{\ell}})\in T_{\mathbf{x}}(\mathbb{R}_{0}^{n+1})^{\ell}$ and $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in(\mathbb{R}_{0}^{n+1})^{\ell}$. For each $\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}$, the orthogonal decomposition induced by $g$ is denoted by:
$T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}=T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}\oplus T_{\mathbf{y}}^{\perp}(\mathcal{E}_{P}^{n})^{\ell},\quad X_{\mathbf{y}}=X_{\mathbf{y}}^{\parallel}+X_{\mathbf{y}}^{\perp},$
@@
$T_{\mathbf{y}}^{\perp}(\mathcal{E}_{P}^{n})^{\ell}=\{X_{\mathbf{y}}\in T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}\mid\langle X_{\mathbf{y}},Y_{\mathbf{y}}\rangle_{g(\mathbf{y})}=0,~\forall~Y_{\mathbf{y}}\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}\}.$
@@
The following lemma gives the explicit expression of the orthogonal projection.
@@
Under the previous conditions, the orthogonal projection is given by $\boldsymbol{\pi}^{\parallel}=T\boldsymbol{\pi}_{P}|_{(\mathcal{E}_{P}^{n})^{\ell}}$.
@@
It is enough to prove that, for each $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in(\mathcal{E}_{P}^{n})^{\ell}$ and $X_{\mathbf{y}}=(X_{{\rm y}_{1}},\ldots,X_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}$, we have that $X_{\mathbf{y}}-T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{\mathbf{y}}\in T_{\mathbf{y}}^{\top}(\mathcal{E}_{P}^{n})^{\ell}$, i.e., that $\langle X_{\mathbf{y}}-T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{\mathbf{y}},Y_{\mathbf{y}}\rangle_{g(\mathbf{y})}=0$ for each $Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. By using Eq.˜2 and Eq.˜7, this latter condition is clearly satisfied:
@@
where we have used Remark˜2.1. ∎
@@ 3.2. Gradient vector field
Let us show that the transformer dynamics is a gradient vector field on the manifold $(\mathcal{E}_{P}^{n})^{\ell}$ equipped with the Riemannian metric $g_{P}=\imath_{P}^{*}g$. For simplicity, we introduce the following vector fields corresponding to Eq.˜6 under ˜3.1 (before and after projecting to the ellipsoid, respectively):
@@
$Y_{P}:(\mathcal{E}_{P}^{n})^{\ell}\to T(\mathcal{E}_{P}^{n})^{\ell},\quad$ $\mathbf{y}\mapsto Y_{P}(\mathbf{y})=\begin{pmatrix}T_{{\rm y}_{1}}\pi_{P}\cdot\left(\sum_{j=1}^{\ell}\alpha_{1j}(\mathbf{y}){\rm y}_{j}\right)\\ \vdots\\ T_{{\rm y}_{\ell}}\pi_{P}\cdot\left(\sum_{j=1}^{\ell}\alpha_{\ell j}(\mathbf{y}){\rm y}_{j}\right)\end{pmatrix}$.
Note that $Y_{P}(\mathbf{y})=T_{\mathbf{y}}\boldsymbol{\pi}_{P}\cdot X_{P}(\mathbf{y})$. Now we show that $X_{P}$ is a gradient field with the metric $g$.
@@
From this, Eq.˜1 and Eq.˜7, we conclude:
${\rm grad}_{g}\,V(\mathbf{x})$ $=-\begin{pmatrix}Z_{1}^{-1}(\mathbf{x})P^{-1}&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&Z_{\ell}^{-1}(\mathbf{x})P^{-1}\end{pmatrix}\begin{pmatrix}\sum_{j=1}^{\ell}\exp({\rm x}_{1}^{\top}P{\rm x}_{j})P{\rm x}_{j}\\ \vdots\\ \sum_{i=j}^{\ell}\exp({\rm x}_{\ell}^{\top}P{\rm x}_{j})P{\rm x}_{j}\end{pmatrix}$
$=-X_{P}(\mathbf{x}).$
@@
where we used Lemma˜3.1 and the equality $T_{\mathbf{y}}\imath_{P}\cdot Y_{P}(\mathbf{y})=Y_{P}(\mathbf{y})$, which follows from regarding $Y_{P}(\mathbf{y})$ both as an element of $T_{\mathbf{y}}(\mathbb{R}_{0}^{n+1})^{\ell}$ and $T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. ∎
@@ 3.3. Stability analysis
Having established that Eq.˜6, under ˜3.1, is a gradient vector field, it is natural to use the potential $V_{P}:(\mathcal{E}_{P}^{n})^{\ell}\to\mathbb{R}$ as a Lyapunov function to study the asymptotic behavior of the tokens. In order to stablish that all trajectories converge to an equilibrium, we first prove that they converge to the zeroes of the gradient field.
Lemma $3.$
The trajectories of Eq.˜6 under ˜3.1 converge to the set:
@@
Let $\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}$ and $Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in T_{\mathbf{y}}(\mathcal{E}_{P}^{n})^{\ell}$. Recall that the formal time derivative (at $t=0$) of the potential $V_{P}$ is the map $\dot{V}_{P}=\mathbf{d}V_{P}(Y_{P}):(\mathcal{E}_{P}^{n})^{\ell}\to\mathbb{R}$. From Theorem˜3.1, we obtain:
@@
If ˜3.1 holds, then every trajectory of Eq.˜6 converges to an equilibrium.
@@
Recall that the potential $V_{P}$ satisfies the Łojasiewicz inequality if $|V_{P}|\leq\lambda|{\rm grad}_{g_{P}}\,V_{P}|_{h}$ for some $\lambda>0$. A sufficient condition for the Łojasiewicz inequality to hold is that $((\mathcal{E}_{P}^{n})^{\ell},g_{P}=\imath^{*}g)$ is a real analytic Riemannian manifold and the potential is real analytic, i.e., $V_{P}\in C^{\omega}((\mathcal{E}_{P}^{n})^{\ell})$. It is clear that these two conditions are satisfied since $(\mathcal{E}_{P}^{n})^{\ell}$ is a real analytic submanifold of $\mathbb{R}^{n+1}$ and $Z_{i},\alpha_{ij}\in C^{\omega}((\mathbb{R}_{0}^{n+1})^{\ell}),\mathbb{R}^{+})$ for each $1\leq i,j\leq\ell$, which ensures that both the Riemannian metric $g_{P}$ and the potential $V_{P}$ are real analytic.
On the other hand, $(\mathcal{E}_{P}^{n})^{\ell}\subset(\mathbb{R}_{0}^{n+1})^{\ell}$ is compact, whence the set of $\omega$-limit points of Eq.˜6 is non-empty. The Łojasiewicz inequality thus ensures that every trajectory converges to a point $\mathbf{y}\in(\mathcal{E}_{P}^{n})^{\ell}$. From Lemma˜3.3, we know that $\mathbf{y}$ is an equilibrium of Eq.˜6 since $\dot{\mathbf{y}}=-{\rm grad}_{g_{P}}\,V_{P}(\mathbf{y})=0$. ∎
@@
4. Full self-attention matrix
In this section we consider a particular case of the model Eq.˜6 described by the following assumptions.
Hypothesis 4.1.
For each head $1\leq\eta\leq h$, we have:
  1. (1)
    $U_{\eta}(t)=\mathbb{I}_{n+1}$,
  2. (2)
    $P_{\eta}(t)$ is bounded, i.e., $\sup_{t\in\mathbb{R}_{0}^{+}}\|P_{\eta}(t)\|<\infty$, and
  3. (3)
    $P_{\eta}(t)$ is uniformly continuous on $\mathbb{R}_{0}^{+}$.
In the proof of Theorem˜4.1 below, a nonsmooth candidate for Lyapunov function will be introduced. In order to handle this situation, we briefly recall how to compute the Dini derivative of a function defined through a maximum (cf. §2.3 of [31]).
@@ 4. Full self-attention matrix
As a particular case, let $\{f_{i}:\mathbb{R}_{0}^{+}\to\mathbb{R}\mid i\in I\}$ be a family of continuously differentiable functions, and consider its maximum:
$f:\mathbb{R}_{0}^{+}\to\mathbb{R},\quad t\mapsto f(t)=\max_{i\in I}f_{i}(t).$
For each $t\in\mathbb{R}_{0}^{+}$, the upper Dini derivative of $f$ is computed according to Danskin’s theorem (cf. Lemma 2.2 of [31]):
@@
where $\mathcal{I}(t)=\{i\in I\mid f(t)=f_{i}(t)\}$. In addition, let us prove the following two lemmas.
@@
If there exists $b>0$ such that $\max_{1\leq\eta\leq h}\sup_{t\in\mathbb{R}_{0}^{+}}\|P_{\eta}(t)\|\leq b$, then there exist $c_{1},c_{2}>0$ such that $c_{1}\leq\alpha_{ij}^{\eta}(t,\mathbf{y})\leq c_{2}$ for each $t\in\mathbb{R}_{0}^{+}$, $\mathbf{y}\in(\mathcal{E}_{W}^{n})^{\ell}$, $1\leq\eta\leq h$ and $1\leq i,j\leq\ell$.
@@
By compactness of $\mathcal{E}_{W}^{n}$, there exists $K>0$ such that $|{\rm y}|\leq K$ for each ${\rm y}\in\mathcal{E}_{W}^{n}$. Therefore, we have $|{\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j}|\leq K^{2}\|P_{\eta}(t)\|\leq K^{2}b$ and, thus, $\exp(-K^{2}b)\leq\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j})\leq\exp(K^{2}b)$. Moreover, we also have:
$\exp(-K^{2}b)\leq\sum_{k=1}^{\ell}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{k})\leq\ell\exp(K^{2}b),$
which leads to:
$\alpha_{ij}^{\eta}(t,\mathbf{y})$ $=\frac{\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{j})}{\sqrt{n+1}\sum_{k=1}^{i}\exp({\rm y}_{i}^{\top}P_{\eta}(t){\rm y}_{k})}$
$\geq\frac{\exp(-K^{2}b)}{\sqrt{n+1}\ell\exp(K^{2}b)}\geq\frac{1}{\sqrt{n+1}\ell\exp(2K^{2}b)}.$
@@
Let $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell}):\mathbb{R}_{0}^{+}\to\mathcal{E}_{W}^{n}$ be a solution of Eq.˜6 under ˜4.1, and $v\in\mathcal{E}_{W}^{n}$. If there exists $b>0$ such that $\sup_{t\in\mathbb{R}_{0}^{+}}\|P_{\eta}(t)\|\leq b$ and $P_{\eta}$ is uniformly continuous on $\mathbb{R}_{0}^{+}$ for each $1\leq\eta\leq h$, then the following functions:
$f_{i}:\mathbb{R}_{0}^{+}\to\mathbb{R},\quad t\mapsto f_{i}(t)=\sum_{\eta=1}^{h}\sum_{j=1}^{\ell}\alpha_{ij}^{\eta}(t,\mathbf{y}(t))\left(v^{\top}{\rm y}_{j}(t)-{\rm y}_{i}(t)^{\top}W{\rm y}_{j}(t)v^{\top}{\rm y}_{i}(t)\right),$
@@
$g_{j}:\mathbb{R}_{0}^{+}\to\mathbb{R},\quad$ $t\mapsto g_{j}(t)=v^{\top}{\rm y}_{j}(t)-{\rm y}_{i}(t)^{\top}W{\rm y}_{j}(t)v^{\top}{\rm y}_{i}(t),$
$g_{j}^{\eta}:\mathbb{R}_{0}^{+}\to\mathbb{R},\quad$ $t\mapsto g_{j}^{\eta}(t)=\alpha_{ij}^{\eta}(t,\mathbf{y}(t)),$
for each $1\leq\eta\leq h$ and $1\leq j\leq\ell$, are bounded: $g_{j}$ due to tokens evolving on the ellipsoid, and $g_{j}^{\eta}$ due to Lemma˜4.1. Recall that the addition and multiplication of bounded and uniformly continuous functions results in uniformly continuous functions. Thus, it is enough to prove that $g_{j}$ and $g_{j}^{\eta}$ are uniformly continuous to conclude that $f_{i}$ is uniformly continuous:
@@
    The derivative of $g_{j}$ is bounded on $\mathbb{R}_{0}^{+}$ since the tokens evolve on the ellipsoid and their dynamics is given by Eq.˜6. Note that $\sup_{t\in\mathbb{R}_{0}^{+}}|\dot{\rm y}_{j}(t)|<\infty$ for each $1\leq j\leq\ell$ thanks to Lemma˜4.1. Hence, $g_{j}$ is uniformly continuous on $\mathbb{R}_{0}^{+}$.
@@
    Given that the tokens evolve on the ellipsoid and every $P_{\eta}$ is bounded on $\mathbb{R}_{0}^{+}$, we can ensure the existence of $K>0$ such that:
    $\max_{1\leq\eta\leq h}\sup_{t\in\mathbb{R}_{0}^{+}}|{\rm y}_{i}(t)^{\top}P_{\eta}(t){\rm y}_{j}(t)|\leq K.$
    Moreover, $\exp|_{[-K,K]}:[-K,K]\to[\exp(-K),\exp(K)]$ is uniformly continuous, as it is defined on a compact. Hence, $\exp({\rm y}_{i}(\cdot)P_{\eta}(\cdot){\rm y}_{j}(\cdot)):\mathbb{R}_{0}^{+}\to[\exp(-K),\exp(K)]$ is uniformly continuous, as the composition of uniformly continuous functions is uniformly continuous. In particular, $Z_{i}^{\eta}(\cdot,\mathbf{y}(\cdot)):\mathbb{R}_{0}^{+}\to[\ell\exp(-K),\ell\exp(K)]$ is uniformly continuous. By gathering all, we conclude that $g_{j}^{\eta}$ is uniformly continuous on $\mathbb{R}_{0}^{+}$.
The next result claims attractivity of the consensus set provided that the initial position of the tokens is in some open hemisphere of the ellipsoid.
@@
If ˜4.1 holds, then the consensus set $\mathcal{C}_{\ell}^{+}(v)$ in the product of hemispheres $\mathcal{H}^{+}(v)^{\ell}$, given by:
@@
is attractive for Eq.˜6 with domain of attraction $\mathcal{H}^{+}(v)^{\ell}$.
@@
For each $t\in\mathbb{R}_{0}^{+}$ and $\mathbf{y}=({\rm y}_{1},\ldots,{\rm y}_{\ell})\in\mathcal{H}^{+}(v)^{\ell}$, note that: